OpenStudy (anonymous):

im giving medals for the best answer In ΔABC shown below, point A is at (0, 0), point B is at (x2, 0), point C is at (x1, y1), point D is at (x sub 1/2 , y sub 1/2) and point E is at ( x sub 1+ x sub 2 /2, y sub 1 /2) Prove that segment DE is parallel to segment AB.

OpenStudy (anonymous):

hold on im trying to get the pic

OpenStudy (anonymous):

@hartnn @amistre64 please help me

OpenStudy (amistre64):

parallel lines have the same slope ....

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

what is the slope from A to B?

OpenStudy (amistre64):

and, what is the slope from D to E?

OpenStudy (anonymous):

im not sure

OpenStudy (amistre64):

you might want to use a slope formula then: given 2 points (a,b) and (p,q) the slope between them is defined to be:\[\frac{b-q}{a-p}\]

OpenStudy (amistre64):

or stated another way:\[slope:\frac{\Delta~y~\text{: change in y}}{\Delta~x\text{: change in x}}\]

OpenStudy (anonymous):

ok now how do i do that

OpenStudy (amistre64):

they give yo the points for A and B; and D and E, plug them into the formula and compare the resutls. if they are equal, they are the same slope, and therefore parallel .... if they are not the same then they are not

OpenStudy (anonymous):

no there are no points did u look at the picture

OpenStudy (amistre64):

um, you posted: n ΔABC shown below, point A is at (0, 0), point B is at (x2, 0), point C is at (x1, y1), point D is at (x sub 1/2 , y sub 1/2) point E is at ( x sub 1+ x sub 2 /2, y sub 1 /2)

OpenStudy (amistre64):

notice the A and B have the same y values; y doesnt change: the formula produces a slope of 0/n = 0 notice the D and E have the same y values; y doesnt change: the formula produces a slope of 0/k = 0 since 0 = 0, they have the same slope

OpenStudy (anonymous):

so that would be my answer??

OpenStudy (amistre64):

that is an overview of what the answer will produce; it is up to you to actually put it into a format that is acceptable by whoever is grading it.

OpenStudy (amistre64):

there are a number of ways that we can show that they lines are parallel, so I cant really say that this is the exact method you should employ. But it should give you an idea on how to approach it

OpenStudy (anonymous):

ok thanks but do u also know someone else that can show me another way